OpenStudy (anonymous):

multiply (5+2√6)(2+√6)

OpenStudy (saifoo.khan):

Using FOIL, \[22 + 9 \sqrt6\]

OpenStudy (tommo_lcfc):

\[(5+2√6)(2+√6) = 10 + 4\sqrt{6} + 5\sqrt{6} + 2\sqrt{36} = 10 + 4\sqrt{6} + 5\sqrt{6} + 12 = 22 + 9\sqrt{6}\]